# Unveiling Over-Memorization in Finetuning LLMs for Reasoning Tasks

### v1

@@ Abstract
The pretrained large language models (LLMs) are finetuned with labeled data for better instruction following ability and alignment with human values. In this paper, we study the learning dynamics of LLM finetuning on reasoning tasks and reveal the uncovered over-memorization phenomenon during a specific stage of LLM finetuning. At this stage, the LLMs have excessively memorized training data and exhibit high test perplexity while maintaining good test accuracy. We investigate the conditions that lead to LLM over-memorization and find that training epochs and large learning rates contribute to this issue. Although models with over-memorization demonstrate comparable test accuracy to normal models, they suffer from reduced robustness, poor out-of-distribution generalization, and decreased generation diversity. Our experiments unveil the over-memorization to be broadly applicable across different tasks, models, and finetuning methods. Our research highlights that overparameterized, extensively finetuned LLMs exhibit unique learning dynamics distinct from traditional machine learning models. Based on our observations of over-memorization, we provide recommendations on checkpoint and learning rate selection during finetuning.
@@ 1 Introduction
In this work, we analyze the learning dynamics of LLM finetuning on reasoning tasks and argue that such heuristic practices may be suboptimal. Specifically, we focus on the study of the over-memorization phenomenon, where we define this as a state where the model has excessively memorized training data, leading to high test perplexity while still counter-intuitively maintaining good test accuracy. We further investigate (1) the conditions inducing the over-memorization phenomenon and (2) its detrimental impacts. Experiments reveal key triggers: higher learning rates accelerate over-memorization onset, while lower learning rates also induce this state given sufficient training epochs, implicating prolonged training duration itself as a fundamental factor regardless of the learning rate. The over-memorization phenomenon is broadly applicable in various tasks, models, as well as finetuning methods like fully finetuning and LoRA-based methods. Although achieving good performance on in-domain benchmarks, over-memorized LLMs tend to exhibit reduced robustness and generalization. Additionally, their abilities to calibrate responses and generate diverse outputs are compromised. These findings suggest that placing too much emphasis on benchmark accuracy can be risky. It may result in selecting models that rely on memorization rather than demonstrating strong generalization capabilities essential for real-world applications. We propose that the characteristic rise in perplexity during over-memorization reflects the model becoming ‘overconfident’ and ‘stubborn’—over-relying on rigid, memorized pathways instead of flexibly exploring alternatives on new data. Recognizing these adverse effects and the phenomenon’s dynamics, we identify the problem of the conventional checkpoint selection methods and propose our suggestions based on the observations in our over-memorization experiments.
@@ 2 Related Work
Deep neural networks often possess more learnable parameters than training samples, enabling them to simply memorize the data rather than converge to generalizable solutions (Novak et al., 2018). Understanding the phenomenon of memorization in models is therefore crucial for improving generalization performance (Brown et al., 2021; Feldman, 2020; Feldman & Zhang, 2020a). Previous studies have explored various definitions of memorization in deep neural networks (Carlini et al., 2019, 2021a; Feldman & Zhang, 2020b; Zhang et al., 2023). One widely accepted definition is based on differential privacy (Dwork et al., 2006), which formalizes the principle that the removal of any single example from the training set should not substantially alter the resulting model. Based on this definition, some works link stronger memorization during training to reduced generalization ability (Bousquet & Elisseeff, 2000), whereas others emphasize its importance for handling long-tail distributions (Feldman, 2020). In language models, memorization typically refers to generating outputs that resemble specific training examples (Carlini et al., 2021b; Inan et al., 2021; Carlini et al., 2023; Tirumala et al., 2022; Hans et al., 2024; Kang et al., 2025). However, some of these works primarily focus on privacy and copyright concerns (Carlini et al., 2021b; Inan et al., 2021; Hans et al., 2024), while others analyze the dynamics of memorization during the pretraining stage (Carlini et al., 2023; Tirumala et al., 2022). The work most similar to ours is Kang et al. (2025), which also investigates memorization during LLM finetuning and its connections to generalization. Kang et al. (2025) specifically analyzes how the finetuned model’s generalization behavior is characterized by its pre-memorization training accuracy. In contrast, to the best of our knowledge, we are the first to uncover the phenomenon of over-memorization and investigate its impact on the model’s generalization behavior.
@@ 3 The Over-Memorization Phenomenon
Inspired by the aforementioned preliminary experiment observation, we study the “over-memorization” phenomenon, where the model has excessively memorized training data, while still counter-intuitively maintaining good test accuracy. [^1] The over-memorization emerges as a phase during LLM finetuning. After more extensive finetuning, the process could transition into a different phrase ‘over-fitting’, where the test accuracy degrades significantly. We focus on the analyses of the over-memorization phenomenon on reasoning tasks like mathematical reasoning. In these tasks, LLMs are finetuned to produce a solution that contains both steps and an answer. While each question is associated with a unique correct answer, the reference output typically illustrates only one possible way to arrive at it. The over-memorized LLMs are likely to generate a reasoning path from the training data and assign high perplexity to other valid paths, while the well-finetuned LLM may generate the correct answer but follow a different reasoning path. Figure 1(right) presents an example that the over-memorized LLM assigns disproportionately low probabilities to alternative tokens or sequences that may still be valid. We further study the over-memorization phenomenon on general tasks in Section 6.1.
@@ 4 When Does The Model Over-Memorize?
In this section, we investigate the factors that affect the over-memorization of LLMs, such as learning rates, training time, and finetuning methods. Our findings indicate that higher learning rates cause the model to enter over-memorization earlier, while lower learning rates also lead to over-memorization after more lengthy training. Furthermore, although different finetuning methods exhibit varying levels of adaptation to the learning rates, they all lead to the model’s over-memorization.
@@ 4.1 Impact of Learning Rates
We finetune the LLM using the LoRA method across five different learning rates and examine their performance on the MetaTrain, MetaTest, and GSM8K data set (see Appendix B.1 for dataset details). As shown in Figure 2, the perplexity on MetaTrain keeps declining, and the model demonstrates varying memorization levels across different learning rates. For instance, a small learning rate (e.g., 2e-6) fails to memorize all data even after 10 epochs, achieving an accuracy below 70%. A learning rate of 2e-4 accelerates memorization compared to 5e-5, suggesting that larger learning rates tend to result in faster learning and memorization of the training data.
@@ 4.3 Impact of Finetuning Methods
The results are shown in Figure 3. Our experiments reveal that over-memorization occurs not only in LoRA but also across various finetuning methods. Specifically, under different finetuning methods, larger learning rates are more likely to lead to over-memorization. Fully finetuning is observed to exhibit higher perplexity compared to LoRA-based methods. Notably, over-memorization in full finetuning not only emerges in lengthy training. As shown in Figure 3(a)-(b), the test perplexity in full finetuning already begins to increase around epoch 3, earlier than in other methods. This suggests that full finetuning may be more prone to early over-memorization, which becomes more prominent as training continues. We attribute this to the fact that fully finetuning updates a larger number of parameters. To further validate this assumption, we design LoRA+, a variant of LoRA. The only difference is that the attention output linear projector and feed-forward gate weight are also finetuned. Under the same settings, LoRA+ exhibits higher perplexity than LoRA, confirming that increasing the number of finetuned parameters leads to higher perplexity. More comprehensive results across a wider range of learning rates, including training, validation, and test performance, are provided in Appendix D.
@@ 5 Behavioral Analyses
Over-memorized models often achieve high test accuracy and seem effective. However, these models typically exhibit higher perplexity on the test set, raising the question of whether they are truly performing well. To address this, we evaluate the model from multiple perspectives, including robustness, OOD performance, diversity, privacy risk, and calibration. To analyze this, we select the LoRA checkpoints from epoch 3 and epoch 10 with a learning rate of 2e-4, representing the normal model and the over-memorized model, respectively. Both models achieve similar performance on the GSM8K test set (75.6 vs. 76.4 for the normal and over-memorized model, respectively), while the test perplexity of the over-memorized model exceeds that of the normal model by 4.52 times. Please refer to Appendix C.1 and Appendix C.2 for details of the evaluations on model calibration and privacy risk.
@@ 5.1 Performance on OOD Dataset
Models ID Test Set OOD Test Set
MetaTest GSM8K MATH GSM-Hard SVAMP TabMWP ASDiv MAWPS Avg.
Normal model 84.5 75.6 28.8 32.9 77.5 67.0 81.2 89.2 69.6
Over-memorized model 92.9 76.4 28.9 31.6 75.4 62.0 79.6 88.7 67.5
In Section 4, we previously observed that the over-memorized model generalizes well on in-distribution (ID) test sets. A natural question that arises is how it performs on out-of-distribution (OOD) test sets. In this section, we explore this by comparing normal and over-memorized models on the OOD mathematical benchmarks, including GSM-Hard (Gao et al., 2023), SVAMP (Patel et al., 2021), TabMWP (Lu et al., 2023), ASDiv (Miao et al., 2020), and MAWPS (Koncel-Kedziorski et al., 2016). We also include results on ID test sets for reference. In addition to MetaTest and GSM8K discussed in Section 3, we add results on MATH (Hendrycks et al., 2021). Since MATH is one of the seed datasets used to synthesize MetaMathQA, it is also considered an ID test set.
As shown in Table 1, while the over-memorized model performs comparably or slightly better than the normal model on the ID test sets, its performance significantly lags behind the normal model on the OOD test sets by an average of 2.1 points. The over-memorized model generalizes well on ID data samples. However, its generalization ability deteriorates significantly when encountering OOD samples. Furthermore, among all the ID test sets, the over-memorized model has better performance on MetaTest compared to the other two ID sets. This is because the distribution of MetaTest is more similar to the training set compared with GSM8K and MATH. Over-memorized models exhibit reduced generalization ability as the distribution gap between the test set and training set increases. In summary, over-memorized models suffer from generalization problems and have poor performance on OOD data, like overfitting models. However, the generalization defects of overfitting models are more pronounced, as they exhibit poor performance even on ID samples.
@@ 5.3 Diversity of Generations
Inference-time techniques that generate multiple outputs (Welleck et al., 2024; Snell et al., 2025), such as Best-of-N (BoN) sampling (Charniak & Johnson, 2005; Stiennon et al., 2020), can enhance LLM reasoning. The effectiveness of these methods often correlates with output diversity, as a wider exploration of the solution space typically yields better results (Chow et al., 2025). This section investigates how such diversity differs between conventionally trained and over-memorized models. To this end, we compare these model types by generating $N$ outputs per query (Figure 4) and evaluate them using two metrics: BoN accuracy, where the optimal sample from $N$ candidates is identified by the Skywork-o1-Open-PRM-Qwen-2.5-7B process reward model (o1 Team, 2024), and Pass@N accuracy, which measures if at least one correct answer is present among the $N$ samples (Chen et al., 2021b).
@@
The results in Figure 4, using a sampling temperature of 1.0, demonstrate that the conventionally trained model consistently surpasses the over-memorized model in both BoN and Pass@N accuracy for $N\geq 3$. To ascertain whether this disparity might be an artifact of an inadequately low temperature for the over-memorized model, its BoN performance was further assessed across temperatures from 1.0 to 2.0. This analysis confirmed that its performance persistently lagged behind the conventionally trained model, especially for larger $N$. These findings suggest that over-memorized models tend to generate less diverse outputs, thereby diminishing the performance gains achievable through repeated sampling techniques for reasoning tasks. This trend is further corroborated by additional lexical and semantic diversity metrics reported in Appendix C.3.
6 Is Over-Memorization Broadly Applicable?
Table 3: Over-memorization across diverse tasks.
Task Metrics Epoch 0 Epoch 2 Epoch 4 Epoch 6 Epoch 8 Epoch 10
GPQA PPL 4.80 4.51 7.47 14.77 30.08 40.53
Accuracy (%) 5.41 21.62 22.30 25.00 22.97 25.00
HumanEval PPL 1.66 1.82 2.13 2.58 3.26 4.08
Accuracy (%) 37.80 46.95 48.17 50.61 48.17 50.61
6.1 Over-Memorization on Diverse Tasks and Diverse Models
Diverse Task
To examine the broader applicability of the over-memorization phenomenon beyond mathematical reasoning, we broaden our study to include code generation, scientific question answering, and open-ended text generation (Appendix D). Across all tasks, we finetune LLaMA-3.1-8B using LoRA (details in Appendix Table 5). For code generation, models are trained on CodeFeedback 10K (Zheng et al., 2025) and evaluated on HumanEval (Chen et al., 2021a). Scientific QA uses the GPQA benchmark (Rein et al., 2024), with 300 training examples and the remainder for evaluation, leveraging gold reasoning chains for perplexity computation.
Diverse Model
(a) (a) Gemma PPL
(b) (a) Gemma Accuracy
(c) (a) Mistral PPL
(d) (a) Mistral Accuracy
6.2 Scaling Analysis
Effect of Model Size
(a) (a) 2e-6
(b) (b) 2e-5
(c) (c) 2e-4
(d) (d) 5e-4
We further investigate the effect of model scale on the over-memorization phenomenon using Qwen2.5 models (Team, 2024b) of varying sizes (0.5B, 1.5B, and 3B), with all finetuning performed via LoRA, consistent with our main experimental setup. Figure 6 reveals several key trends: (1) Larger models generally yield better test performance. (2) Over-memorization is more pronounced in larger models, as indicated by substantially higher test perplexity in the overtraining regime (Figures 6(c)–(d)). (3) Larger models exhibit stronger initial generalization, achieving lower test perplexity early in training. However, their greater capacity for memorization (Tirumala et al., 2022) leads to a faster increase in test perplexity over time, making them more susceptible to over-memorization (Figures 6(c)–(d)). (4) Larger models favor smaller optimal learning rates, as shown in Figure 14 (Appendix D). This may be due to the aggregate effect of conservative updates across a large number of parameters, enabling effective optimization even with small per-parameter steps.
Effect of Data Size
The results in Figure 8 (see Appendix D.1) reveal several key insights: (1) The over-memorization pattern—characterized by rising test perplexity despite stable accuracy—persists across all dataset sizes. (2) In line with prior work (Zhang et al., 2024a), larger datasets generally yield better test performance. (3) Most notably, we observe a systematic interaction between dataset size and optimal learning rate: larger datasets favor smaller learning rates. For instance, the optimal learning rate shifts from 2e-4 on the 5K subset to 2e-5 on the 100K subset, with the 30K subset performing well under both. This behavior likely stems from the increased number of optimization steps induced by larger datasets under fixed training epochs and batch size, making smaller learning rates sufficient to traverse the solution space.
7 Discussions
7.1 Recipes of Checkpoint Selection

### v2

@@ Abstract
The pretrained large language models (LLMs) are finetuned with labeled data for better instruction following ability and alignment with human values. In this paper, we study the learning dynamics of LLM finetuning on reasoning tasks and reveal the uncovered over-memorization phenomenon during a specific stage of LLM finetuning. At this stage, the LLMs have excessively memorized training data and exhibit high test perplexity while maintaining good test accuracy. We explore the conditions that contribute to over-memorization and discover that this issue is prevalent across various tasks, models, and fine-tuning methods, with prolonged training and large learning rates exacerbating the problem. Although models with over-memorization demonstrate comparable test accuracy to normal models, they suffer from reduced robustness, poor out-of-distribution generalization, and decreased generation diversity. In light of our findings on over-memorization, we offer recommendations for checkpoint selection and propose techniques such as checkpoint merging and memorization-aware reweighting to mitigate this effect.
@@ 1 Introduction
In this work, we analyze the learning dynamics of LLM finetuning on reasoning tasks and argue that such heuristic practices may be suboptimal. Specifically, we focus on the study of the over-memorization phenomenon, where we define this as a state where the model has excessively memorized training data, leading to high test perplexity while still counter-intuitively maintaining good test accuracy. We further investigate (1) the conditions inducing the over-memorization phenomenon and (2) its detrimental impacts. Experiments reveal key triggers: higher learning rates accelerate over-memorization onset, while lower learning rates also induce this state given sufficient training epochs, implicating prolonged training duration itself as a fundamental factor regardless of the learning rate. The over-memorization phenomenon is broadly applicable in various tasks, models, as well as finetuning methods like fully finetuning and LoRA-based methods. Although achieving good performance on in-domain benchmarks, over-memorized LLMs tend to exhibit reduced robustness and generalization. Additionally, their abilities to calibrate responses and generate diverse outputs are compromised.
These findings suggest that placing too much emphasis on benchmark accuracy can be risky. It may result in selecting models that rely on memorization rather than demonstrating strong generalization capabilities essential for real-world applications. We propose that the characteristic rise in perplexity during over-memorization reflects the model becoming ‘overconfident’ and ‘stubborn’—over-relying on rigid, memorized pathways instead of flexibly exploring alternatives on new data. Recognizing these adverse effects and the phenomenon’s dynamics, we identify the problem of the conventional checkpoint selection methods and propose our suggestions based on the observations in our over-memorization experiments. Additionally, we explore two lightweight mitigation strategies, checkpoint merging and memory-aware reweighting loss, which effectively alleviate over-memorization.
@@ 2 Related Work
Deep neural networks often possess more learnable parameters than training samples, enabling them to simply memorize the data rather than converge to generalizable solutions (Novak et al., 2018). Understanding the phenomenon of memorization in models is therefore crucial for improving generalization performance (Brown et al., 2021; Feldman, 2020; Feldman & Zhang, 2020a). Previous studies have explored various definitions of memorization in deep neural networks (Carlini et al., 2019; 2021a; Feldman & Zhang, 2020b; Zhang et al., 2023). One widely accepted definition is based on differential privacy (Dwork et al., 2006), which formalizes the principle that the removal of any single example from the training set should not substantially alter the resulting model. Based on this definition, some works link stronger memorization during training to reduced generalization ability (Bousquet & Elisseeff, 2000), whereas others emphasize its importance for handling long-tail distributions (Feldman, 2020). In language models, memorization typically refers to generating outputs that resemble specific training examples (Carlini et al., 2021b; Inan et al., 2021; Carlini et al., 2023; Tirumala et al., 2022; Hans et al., 2024). However, some of these works primarily focus on privacy and copyright concerns (Carlini et al., 2021b; Inan et al., 2021; Hans et al., 2024), while others analyze the dynamics of memorization during the pretraining stage (Carlini et al., 2023; Tirumala et al., 2022). The work most similar to ours is Kang et al. (2025), which also investigates memorization during LLM finetuning and specifically analyzes how the finetuned model’s generalization behavior is characterized by its pre-memorization training accuracy. In contrast, to the best of our knowledge, we are the first to uncover the phenomenon of over-memorization and investigate its impact on the model’s generalization behavior.
@@ 3 The Over-Memorization Phenomenon
Inspired by the aforementioned preliminary experiment observation, we study the “over-memorization” phenomenon, where the model has excessively memorized training data, while still counter-intuitively maintaining good test accuracy. We focus on the analyses of the over-memorization phenomenon on reasoning tasks like mathematical reasoning. In these tasks, LLMs are finetuned to produce a solution that contains both steps and an answer. While each question is associated with a unique correct answer, the reference output typically illustrates only one possible way to arrive at it. The over-memorized LLMs are likely to generate a reasoning path from the training data and assign high perplexity to other valid paths, while the well-finetuned LLM may generate the correct answer but follow a different reasoning path. Figure 1(right) presents an example that the over-memorized LLM assigns disproportionately low probabilities to alternative tokens or sequences that may still be valid. We further study the over-memorization phenomenon on general tasks in Section 4.4. A brief mechanistic explanation is provided in Appendix C.
@@ 4 When Does The Model Over-Memorize?
In this section, we investigate the factors that affect the over-memorization of LLMs, such as learning rates, training time, and finetuning methods. We further extend our analyses to different models and tasks, and provide scaling experiments on dataset size and model scale (Appendix E.1). Our findings indicate that higher learning rates cause the model to enter over-memorization earlier, while lower learning rates also lead to over-memorization after more lengthy training. Furthermore, although different finetuning methods exhibit varying levels of adaptation to the learning rates, they all lead to the model’s over-memorization. By default, we finetune LLaMA-3.1- 8B on the MetaMathQA 100K dataset and examine their performance on the MetaTrain, MetaTest, and GSM8K dataset (see Appendix B.1 for dataset details) unless otherwise specified.
@@ 4.1 Impact of Learning Rates
We finetune the LLM using the LoRA method across five different learning rates. As shown in Figure 2, the perplexity on MetaTrain keeps declining, and the model demonstrates varying memorization levels across different learning rates. For instance, a small learning rate (e.g., 2e-6) fails to memorize all data even after 10 epochs, achieving an accuracy below 70%. A learning rate of 2e-4 accelerates memorization compared to 5e-5, suggesting that larger learning rates tend to result in faster learning and memorization of the training data.
@@ 4.3 Impact of Finetuning Methods
The results are shown in Figure 3. Our experiments reveal that over-memorization occurs not only in LoRA but also across various finetuning methods. Specifically, under different finetuning methods, larger learning rates are more likely to lead to over-memorization. Fully finetuning is observed to exhibit higher perplexity compared to LoRA-based methods. Notably, over-memorization in full finetuning not only emerges in lengthy training. As shown in Figure 3(a)-(b), the test perplexity in full finetuning already begins to increase around epoch 3, earlier than in other methods. This suggests that full finetuning may be more prone to early over-memorization, which becomes more prominent as training continues. We attribute this to the fact that fully finetuning updates a larger number of parameters. To further validate this assumption, we design LoRA+, a variant of LoRA. The only difference is that the attention output linear projector and feed-forward gate weight are also finetuned. Under the same settings, LoRA+ exhibits higher perplexity than LoRA, confirming that increasing the number of finetuned parameters leads to higher perplexity. More comprehensive results across a wider range of learning rates, including training, validation, and test performance, are provided in Appendix E.
4.4 Over-Memorization on Diverse Tasks and Diverse Models
Table 1: Over-memorization across diverse tasks.
Task Metrics Epoch 0 Epoch 2 Epoch 4 Epoch 6 Epoch 8 Epoch 10
GPQA PPL 4.80 4.51 7.47 14.77 30.08 40.53
Accuracy (%) 5.41 21.62 22.30 25.00 22.97 25.00
HumanEval PPL 1.66 1.82 2.13 2.58 3.26 4.08
Accuracy (%) 37.80 46.95 48.17 50.61 48.17 50.61
Diverse Task
To examine the broader applicability of the over-memorization phenomenon beyond mathematical reasoning, we broaden our study to include code generation, scientific question answering, and open-ended text generation (Appendix E). Across all tasks, we finetune LLaMA-3.1-8B using LoRA (details in Appendix Table 6). For code generation, models are trained on CodeFeedback 10K (Zheng et al., 2025) and evaluated on HumanEval (Chen et al., 2021a). Scientific QA uses the GPQA benchmark (Rein et al., 2024), with 300 training examples and the remainder for evaluation, leveraging gold reasoning chains for perplexity computation.
Diverse Model
(a) (a) Gemma PPL
(b) (b) Gemma Accuracy
(c) (c) Mistral PPL
(d) (d) Mistral Accuracy
@@ 5 Behavioral Analyses
Over-memorized models often achieve high test accuracy and seem effective. However, these models typically exhibit higher perplexity on the test set, raising the question of whether they are truly performing well. To address this, we evaluate the model from multiple perspectives, including robustness, OOD performance, diversity, privacy risk, and calibration. To analyze this, we select the LoRA checkpoints from epoch 3 and epoch 10 with a learning rate of 2e-4, representing the normal model and the over-memorized model, respectively. Both models achieve similar performance on the GSM8K test set (75.6 vs. 76.4 for the normal and over-memorized model, respectively), while the test perplexity of the over-memorized model exceeds that of the normal model by 4.52 times. Please refer to Appendix D.1 and Appendix D.2 for details of the evaluations on model calibration and privacy risk.
@@ 5.1 Performance on OOD Dataset
Methods ID Test Set OOD Test Set
GSM8K MATH SVAMP ASDiv MAWPS TabMWP Minerva MMLU_STEM Avg.
Normal model 75.6 28.9 77.1 81.7 89.7 67.1 27.4 17.9 60.2
Over-memorized model 76.4 28.5 75.3 79.1 89.6 63.6 29.2 12.0 58.1
In Section 4, we previously observed that the over-memorized model generalizes well on in-distribution (ID) test sets. A natural question that arises is how it performs on out-of-distribution (OOD) test sets. In this section, we explore this by comparing normal and over-memorized models on the OOD mathematical benchmarks, including SVAMP (Patel et al., 2021), ASDiv (Miao et al., 2020), MAWPS (Koncel-Kedziorski et al., 2016), TabMWP (Lu et al., 2023), Minerva_MATH (Lewkowycz et al., 2022), and MMLU_STEM (Hendrycks et al., 2020). We also include results on ID test sets for reference. In addition to GSM8K discussed in Section 3, we add results on MATH (Hendrycks et al., 2021). Since MATH is one of the seed datasets used to synthesize MetaMathQA, it is also considered an ID test set.
As shown in Table 2, the over-memorized model performs comparably to the normal model on in-domain (ID) test sets, but falls behind by an average of 2.1 points on out-of-domain (OOD) benchmarks. This suggests that over-memorization does not harm ID generalization, but undermines robustness to distribution shifts. The effect becomes more pronounced as the gap between training and test distributions increases, indicating that over-memorized models are particularly vulnerable to OOD evaluation. In summary, over-memorized models suffer from generalization problems and have poor performance on OOD data, like overfitting models. However, the generalization defects of overfitting models are more pronounced, as they exhibit poor performance even on ID samples.
@@ 5.3 Diversity of Generations
Inference-time techniques that generate multiple outputs (Welleck et al., 2024; Snell et al., 2025), such as Best-of-N (BoN) sampling (Charniak & Johnson, 2005; Stiennon et al., 2020), can enhance LLM reasoning. The effectiveness of these methods often correlates with output diversity, as a wider exploration of the solution space typically yields better results (Chow et al., 2025). To this end, we compare these model types by generating $N$ outputs per query (Figure 5) and evaluate them using two metrics: BoN accuracy, where the optimal sample from $N$ candidates is identified by the Skywork-o1-Open-PRM-Qwen-2.5-7B process reward model (o1 Team, 2024), and Pass@N accuracy, which measures if at least one correct answer is present among the $N$ samples (Chen et al., 2021b).
@@
The results in Figure 5, using a sampling temperature of 1.0, demonstrate that the conventionally trained model consistently surpasses the over-memorized model in both BoN and Pass@N accuracy for $N\geq 3$. To ascertain whether this disparity might be an artifact of an inadequately low temperature for the over-memorized model, its BoN performance was further assessed across temperatures from 1.0 to 2.0. This analysis confirmed that its performance persistently lagged behind the conventionally trained model, especially for larger $N$. These findings suggest that over-memorized models tend to generate less diverse outputs, thereby diminishing the performance gains achievable through repeated sampling techniques for reasoning tasks. This trend is further corroborated by additional lexical and semantic diversity metrics reported in Appendix D.3.
6 Methods for Mitigating Over-Memorization
6.1 Receipt for Checkpoint Selection